i donot know how to slove this problem:
2x-8=10
please, step by step show me how to answer because my teacher CANOT teach

Answer by askmemath(368)

Adding 8 on both sides
2x-8+8 = 10+8
2x+0 = 18
2x = 18
Dividing by 2 throughout
X = 9